# Realm: Recursive Relevance Modeling for LLM-based Document Re-Ranking

### v1

@@ Abstract
Large Language Models (LLMs) have shown strong capabilities in document re-ranking, a key component in modern Information Retrieval (IR) systems. However, existing LLM-based approaches face notable limitations, including ranking uncertainty, unstable top-$k$ recovery, and high token cost due to token-intensive prompting. To effectively address these limitations, we propose Realm, an uncertainty-aware re-ranking framework that models LLM-derived relevance as Gaussian distributions and refines them through recursive Bayesian updates. By explicitly capturing uncertainty and minimizing redundant queries, Realm achieves better rankings more efficiently. Experimental results demonstrate that our Realm surpasses state-of-the-art re-rankers while significantly reducing token usage and latency, promoting it as the next generation re-ranker for modern IR systems.
Realm: Recursive Relevance Modeling for LLM-based Document Re-Ranking
Pinhuan Wang1, Zhiqiu Xia1, Chunhua Liao2, Feiyi Wang3, Hang Liu1, 1 Rutgers, The State University of New Jersey, 2 Lawrence Livermore National Laboratory, 3 Oak Ridge National Laboratory,
@@ 1 Introduction
Document re-ranking is a key component in modern IR systems Zhu et al. (2024). Given a user query, retrieval systems typically begin with a fast but coarse retrieval stage that returns a broad set of potentially relevant documents. However, these initial results are often noisy or only loosely related to the query. Re-ranking addresses this issue by applying a more accurate, context-aware scoring model to refine the order of the candidates and place the most relevant documents at the top Nogueira and Cho (2020). For example, in academic paper searching, an initial retrieval step may return a large number of documents that match surface-level keywords, while missing more in-depth relevance analysis. Without re-ranking, users could suffer from missing important references, thus struggling to construct a complete picture of the research landscape. Recent work Fukuda and Tomiura (2025) shows that integrating multiple retrieval methods, particularly re-ranking the combined results, significantly improves recall, helping users retrieve more comprehensive and relevant literature within a manageable number of results. In a nutshell, re-ranking is crucial to ensure that high-quality, contextually appropriate documents are selected as input for subsequent applications Lewis et al. (2020).
LLMs are redefining document re-ranking by enabling deep semantic and contextual understanding that traditional lexical-based methods fundamentally lack. Traditional re-rankers, such as those based on BM25 scores Robertson et al. (2009) or learning-to-rank models like LambdaMART Burges (2010), rely heavily on sparse features-term overlap, document frequency, or hand-crafted heuristics-which often fail in capturing nuanced relevance, especially in complex or ambiguous queries. In contrast, LLM-based re-rankers treat the query and candidate documents as joint inputs, allowing for fine-grained relevance estimation grounded in deep semantic and contextual comprehension. Recent studies have shown that LLMs, when applied as cross-encoders or guided with task-specific prompting, consistently outperform classical re-rankers across benchmarks Sun et al. (2023). These advancements suggest that LLMs are not just an incremental improvement but a paradigm shift toward unifying retrieval and comprehension within a single, adaptable framework.
@@
Recent research endeavors have fallen short in effectively addressing all three challenges: Pointwise methods Nogueira et al. (2020) are efficient and parallelizable, as they assess each document independently. Some variants Zhuang et al. (2023b) further leverage generation likelihood as a relevance score. However, pointwise approaches fail to model interactions among candidates, making them less effective at resolving uncertainty or producing globally consistent top-$k$ rankings. Listwise methods Sun et al. (2023) enable joint evaluation of multiple candidates in a single query, which helps mitigate ranking inconsistency. Approaches like TourRank Chen et al. (2025) adopt tournament-style aggregation to extend listwise scoring. Despite these benefits, listwise methods still suffer from context length constraints and positional bias Liu et al. (2024b), especially for long candidate sets. Pairwise methods Qin et al. (2024) improve local comparison quality by directly modeling relative preferences between document pairs. Advanced systems like PRP-Graph Luo et al. (2024) further exploit graph structures to aggregate pairwise signals. Nevertheless, the repeated comparisons, however, lead to high token usage and substantial inference latency. Setwise methods Zhuang et al. (2024); Podolak et al. (2025) improve efficiency by evaluating small subsets at a time, but discard fine-grained preference information, such as full relevance logits—thereby under-utilizing the model’s capacity.
@@ 2.1 Related Work
Listwise methods. With the continued expansion of LLM capacity and input window size, listwise ranking—where the model receives a group of candidate documents and directly outputs their relative ordering—has become increasingly feasible. By supporting joint reasoning over multiple candidates within a single inference, this paradigm has motivated a series of methods aimed at better leveraging LLM capabilities for document re-ranking.
@@ 3.1 Realm Framework
Consequently, we introduce a recursive design, that is, we compare each document against the pivot document. If a document is closer to the query than the pivot, we keep it for the next round of calculation. Otherwise, we filter out that document. Moreover, directly relying on the pivot $D_{3}$ to filter out unpromising documents would yield unstable workload reduction. We thus design an effective workload reduction mechanism to cope with this concern, which could derive $D_{4}$ as the final split point ( 4 ). Of note, we also design strategy to select the document with the highest confidence as the pivot (step  1.2 ).
@@ 3.2 Realm’s Relevance Modeling Scheme
In setwise comparison, LLM is prompted to select the most relevant document from a small group (of size $m$) of candidate documents. Our example query is “Given a query, which of the following passages is the most relevant?”. Internally, LLM assigns scalar logits to each option; we extract these as a score vector $\{\ell_{i}\colon D_{i}\in\text{set}\}$, where each $\ell_{i}$ corresponds to the document $D_{i}$. For instance, in step  2.1 of Figure 1, the logits assigned to $(D_{0},D_{1},D_{3})$ are $3.2$, $1.1$, and 0.8-0.8, respectively. Of note, for closed-source LLMs, we can prompt them to output the confidence values and use that as the scores Xia et al. (2025).
@@
where $\sigma$ is the sigmoid function and $T$ is a temperature parameter. This formulation transforms a single multi-document comparison into a set of $\binom{m}{2}$ pairwise probabilities updates.
@@
As illustrated in step  2.2 , the model predicts $P(D_{0}\succ D_{3})=0.73$. In this case, the relevance distribution of $D_{0}$, i.e., $\mathcal{N}(\mu_{0,0},\sigma_{0,0}^{2})$, is updated to $\mathcal{N}(\mu_{0,1},\sigma_{0,1}^{2})$ by interpolating between the two TrueSkill-updated distributions—corresponding to a win $\mathcal{N}(\mu_{win},\sigma_{win}^{2})$ (green dashed curve) and a loss $\mathcal{N}(\mu_{loss},\sigma_{loss}^{2})$ (red dashed curve) against $D_{3}$—weighted by the predicted probability.
@@ 3.3 Pivot-Centric Optimizations
Pivot selection. To determine the global pivot at each recursive step, we select the document with the lowest estimated standard deviation $\sigma$ from the current candidate pool. Intuitively, a lower $\sigma$ indicates higher confidence in the document’s relevance estimate. Using such a document as the pivot improves the stability of the partitioning process. As illustrated in step  1.2 of Figure 1, document $D_{3}$ is selected as the global pivot due to its lowest uncertainty. The remaining documents are then grouped into prompts: $\{D_{0},D_{1}\}$, $\{D_{2},D_{4}\}$, and $\{D_{5}\}$, ensuring that each non-pivot document is compared once against $D_{3}$. This structure enables us to construct a globally consistent relevance preference centered on a high-confidence document.
@@
Table 1: Evaluation on TREC DL 2019 and TREC DCL 2020 datasets: Realm vs TourRank Chen et al. (2025), PRP-Graph Luo et al. (2024), Setwise-Heapsort Zhuang et al. (2024), and Setwise-Insertion Podolak et al. (2025).
@@ 4.1 Experimental Setup
We conduct evaluations on Flan-T5 models Longpre et al. (2023) of three sizes—Flan-T5-Large (770M parameters), Flan-T5-XL (3B), and Flan-T5-XXL (11B)—following recent work on LLM-based re-ranking Qin et al. (2024); Luo et al. (2024); Zhuang et al. (2024); Podolak et al. (2025). To assess the generality of Realm, we additionally evaluate Flan-UL2 (20B) Tay et al. (2023) and LLaMA3 (8B and 70B) Grattafiori et al. (2024). Notably, LLaMA3 follows a decoder-only architecture and differs from Flan-T5 models in both model structure and pretraining objectives.
@@
Datasets and metrics. Experiments are conducted on two widely used benchmarks: TREC Deep Learning 2019 Craswell et al. (2020) and 2020 Craswell et al. (2021). All LLM-based methods re-rank the top $100$ documents retrieved by a BM25 first-stage retriever. Adopting prior work’s evaluation strategy Zhuang et al. (2024); Podolak et al. (2025), we formulate re-ranking as a top-$k$ task, with $k=10$ as the default setting. Effectiveness is measured using the NDCG@10 metric for all datasets. For clarity, all NDCG@10 results are presented as percentages..
@@ 4.2 Overall Evaluation
Table 1 presents a comprehensive comparison of our projects on both the TREC-DL 2019 and 2020 benchmarks. We report NDCG@10 (N@10), Inf. (inference counts, a.k.a., the # of LLM calls), P. tks. (#tokens in prompt), G. tks. (# of generated tokens), and Lat. (latency in seconds).
@@ 4.3 Analysis
Removing any of these components leads to a consistent drop in performance. Disabling uncertainty modeling (w/o modeling) results in a 1.53.31.5-3.3 decrease in NDCG@10 across all models, highlighting the value of Gaussian-based relevance modeling. The absence of recursive reasoning (w/o recursive) also causes noticeable degradation, underscoring the benefit of multi-round refinement. Lastly, disabling pivot-centric optimization (w/o optimization) nearly doubles latency—for example, from $7.4$s to $12.1$s with Flan-T5-XXL—confirming that our pivot selection and partitioning strategy substantially improves efficiency without compromising effectiveness.
@@ 5 Conclusion
Our current design choices are partially constrained by resource and space limitations. While our relevance modeling is, in principle, compatible with a broad range of re-ranking methods, we center our evaluation on the specific framework we designed that achieved the strongest empirical performance. A more comprehensive comparison with alternative designs is left to future work. Secondly, we conduct experiments using open-source models, including the Flan-T5 and LLaMA 3 series. We did not include closed-source models such as GPT-3.5 or GPT-4 accessed via API. Future extensions may consider such models to provide a more complete empirical picture of Realm.
@@ B.1 Detailed Explanation of Datasets
The TREC Deep Learning (DL) 2019 Craswell et al. (2020) and 2020 Craswell et al. (2021) datasets are benchmark collections designed to evaluate document ranking systems in complex information retrieval tasks. Both datasets are based on queries derived from real-world search logs and are built on top of the MS MARCO Nguyen et al. (2016) passage and document corpora. The TREC DL 2019 dataset includes $43$ queries with graded relevance judgments, while the 2020 version expands the test set to $54$ queries. Documents are written in English and come from a large web-scale corpus, with each query typically associated with hundreds to thousands of candidate documents. These datasets emphasize fine-grained relevance estimation and are widely used for evaluating neural re-ranking models and LLM-based retrieval methods. The reported NDCG@10 scores are averaged over the entire dataset with a single run, ensuring stable and reliable evaluation results.

### v2

# Pinhuan Wang1Zhiqiu Xia1Chunhua Liao2Feiyi Wang3Hang Liu1

@@ Abstract
Large Language Models (LLMs) have shown strong capabilities in document re-ranking, a key component in modern Information Retrieval (IR) systems. However, existing LLM-based approaches face notable limitations, including ranking uncertainty, unstable top-$k$ recovery, and high token cost due to token-intensive prompting. To effectively address these limitations, we propose Realm, an uncertainty-aware re-ranking framework that models LLM-derived relevance as Gaussian distributions and refines them through recursive Bayesian updates. By explicitly capturing uncertainty and minimizing redundant queries, Realm achieves better rankings more efficiently. Experimental results demonstrate that our Realm surpasses state-of-the-art re-rankers while significantly reducing token usage and latency, improving NDCG@10 by 0.711.90.7-11.9 and simultaneously reducing the number of LLM inferences by $23.4-84.4\%$, promoting it as the next-generation re-ranker for modern IR systems.
Realm: Recursive Relevance Modeling for LLM-based Document Re-Rankingthanks: Realm is integrated in: https://ipapers.ai/.
Pinhuan Wang1   Zhiqiu Xia1   Chunhua Liao2   Feiyi Wang3   Hang Liu1 1Rutgers, The State University of New Jersey 2Lawrence Livermore National Laboratory  3Oak Ridge National Laboratory {pinhuan.wang, hang.liu}@rutgers.edu
@@ 1 Introduction
Document re-ranking is a key component in modern IR systems Zhu et al. (2024). Given a user query, retrieval systems typically begin with a fast but coarse retrieval stage that returns a broad set of potentially relevant documents. However, these initial results are often noisy or only loosely related to the query. Re-ranking addresses this issue by applying a more accurate, context-aware scoring model to refine the order of the candidates and place the most relevant documents at the top Nogueira and Cho (2020). For example, in academic paper searching, an initial retrieval step may return a large number of documents that match surface-level keywords, while missing more in-depth relevance analysis. Without re-ranking, the most important references could be placed at the bottom, potentially leading researchers to miss those key published findings in their research journey. In a nutshell, re-ranking is crucial to ensure that high-quality, contextually appropriate documents are selected as input for subsequent applications Lewis et al. (2020).
LLMs are redefining document re-ranking by enabling deep semantic and contextual understanding that traditional methods fundamentally lack. Traditional re-rankers, such as those based on BM25 scores Robertson and Zaragoza (2009) or learning-to-rank models like LambdaMART Burges (2010), rely on either simple sparse feature-term overlap, document frequency, or hand-crafted heuristics, which often fail in capturing nuanced relevance, especially in complex or fine-grained queries. In contrast, LLM-based re-rankers treat the query and candidate documents as joint inputs, allowing for more in-depth relevance estimation grounded in deep semantic and contextual comprehension. Recent studies have shown that LLMs, when applied as cross-encoders or guided with task-specific prompting, consistently outperform classical re-rankers across benchmarks Sun et al. (2023). These advancements suggest that LLMs are not just an incremental improvement but a paradigm shift toward unifying retrieval and comprehension within a single, adaptable framework.
@@
Recent research endeavors have fallen short in effectively addressing all three challenges: Pointwise methods Nogueira et al. (2020) are efficient and parallelizable, as they assess each document independently. Some variants Zhuang et al. (2023b) further leverage generation likelihood as a relevance score. However, pointwise approaches fail to model interactions among candidates, making them less effective at resolving uncertainty or producing globally consistent top-$k$ rankings. Listwise methods Sun et al. (2023) enable joint evaluation of multiple candidates in a single query, which helps mitigate ranking inconsistency. Approaches like TourRank Chen et al. (2025) adopt tournament-style aggregation to extend listwise scoring. Despite these benefits, listwise methods still suffer from context length constraints and positional bias Liu et al. (2024b), especially for long candidate sets. Pairwise methods Qin et al. (2024) improve local comparison quality by directly modeling relative preferences between document pairs. Advanced systems like PRP-Graph Luo et al. (2024) further exploit graph structures to aggregate pairwise signals. Nevertheless, the repeated comparisons lead to high token usage and substantial inference latency. Setwise methods Zhuang et al. (2024); Podolak et al. (2025) improve efficiency by evaluating small subsets at a time, but discard fine-grained preference information, such as full relevance logits—thereby under-utilizing the model’s capacity.
@@ 2.1 Related Work
Listwise methods. With the continued expansion of LLM capacity and input window size, listwise ranking, where the model receives a group of candidate documents and directly outputs their relative ordering, has become increasingly feasible. By supporting joint reasoning over multiple candidates within a single inference, this paradigm has motivated a series of methods aimed at better leveraging LLM capabilities for document re-ranking.
@@ 3.1 Realm Framework
Consequently, we introduce a recursive design, that is, we compare each document against the pivot document. If a document is closer to the query than the pivot, we keep it for the next round of calculation. Otherwise, we filter out that document. Moreover, directly relying on the pivot $D_{3}$ to filter out unpromising documents would yield unstable workload reduction. We thus design an effective workload reduction mechanism to cope with this concern, which could derive $D_{4}$ as the final split point ( 4 ). Of note, we also design a strategy to select the document with the highest confidence as the pivot (step  1.2).
@@ 3.2 Realm’s Relevance Modeling Scheme
In setwise comparison, LLM is prompted to select the most relevant document from a small group (of size $m$) of candidate documents. Our example query is "Given a query, which of the following passages is the most relevant?". Internally, LLM assigns scalar logits to each option; we extract these as a score vector $\{\ell_{i}\colon D_{i}\in\text{set}\}$, where each $\ell_{i}$ corresponds to the document $D_{i}$. For instance, in step  2.1 of Figure 1, the logits assigned to $(D_{0},D_{1},D_{3})$ are $3.2$, $1.1$, and 0.8-0.8, respectively. Of note, for closed-source LLMs, we can prompt them to output the confidence values and use that as the scores Xia et al. (2025).
@@
where $\sigma$ is the sigmoid function and $T$ is a temperature parameter. This formulation transforms a single multi-document comparison into a set of $\binom{m}{2}$ pairwise probability updates.
@@
As illustrated in step  2.2, the model predicts $P(D_{0}\succ D_{3})=0.73$. In this case, the relevance distribution of $D_{0}$, i.e., $\mathcal{N}(\mu_{0,0},\sigma_{0,0}^{2})$, is updated to $\mathcal{N}(\mu_{0,1},\sigma_{0,1}^{2})$ by interpolating between the two TrueSkill-updated distributions: a win $\mathcal{N}(\mu_{win},\sigma_{win}^{2})$ (green dashed curve) and a loss $\mathcal{N}(\mu_{loss},\sigma_{loss}^{2})$ (red dashed curve) against $D_{3}$, weighted by the predicted probability.
@@ 3.3 Pivot-Centric Optimizations
Pivot selection. To determine the global pivot at each recursive step, we select the document with the lowest estimated standard deviation $\sigma$ from the current candidate pool. Intuitively, a lower $\sigma$ indicates higher confidence in the document’s relevance estimate. Using such a document as the pivot improves the stability of the partitioning process. Initially, the $\sigma$ may be the same across documents. In this case, we select the document with the median estimated relevance score $\mu$ as the pivot, to avoid pruning too many or too few candidates. As illustrated in step  1.2 of Figure 1, document $D_{3}$ is selected as the global pivot due to its lowest uncertainty. The remaining documents are then grouped into prompts: $\{D_{0},D_{1}\}$, $\{D_{2},D_{4}\}$, and $\{D_{5}\}$, ensuring that each non-pivot document is compared once against $D_{3}$. This structure enables us to construct a globally consistent relevance preference centered on a high-confidence document.
@@
Table 1: Evaluation on TREC DL 2019 and TREC DL 2020 datasets: Realm vs TourRank Chen et al. (2025), PRP-Graph Luo et al. (2024), Setwise-Heapsort Zhuang et al. (2024), and Setwise-Insertion Podolak et al. (2025).
@@ 4.1 Experimental Setup
We conduct evaluations on Flan-T5 models Longpre et al. (2023) of three sizes—Flan-T5-Large (770M parameters), Flan-T5-XL (3B), and Flan-T5-XXL (11B)-following recent work on LLM-based re-ranking Qin et al. (2024); Luo et al. (2024); Zhuang et al. (2024); Podolak et al. (2025). To assess the generality of Realm, we additionally evaluate Flan-UL2 (20B) Tay et al. (2023) and LLaMA3 (8B and 70B) Grattafiori et al. (2024). Notably, LLaMA3 follows a decoder-only architecture and differs from Flan-T5 models in both model structure and pretraining objectives.
@@
Datasets and metrics. Experiments are conducted on widely used benchmarks: TREC Deep Learning 2019 Craswell et al. (2020), 2020 Craswell et al. (2021), and the BEIR benchmark Thakur et al. (2021).
All LLM-based methods re-rank the top $100$ documents retrieved by a BM25 first-stage retriever. Adopting prior work’s evaluation strategy Zhuang et al. (2024); Podolak et al. (2025), we formulate re-ranking as a top-$k$ task, with $k=10$ as the default setting. Effectiveness is measured using the NDCG@10 metric for all datasets. For clarity, all NDCG@10 results are presented as percentages.
@@ 4.2 Overall Evaluation
Table 1 presents a comprehensive comparison of our approach on both the TREC-DL 2019 and 2020 benchmarks. We report NDCG@10 (N@10), Inf. (inference counts, a.k.a., the # of LLM calls), P. tks. (#tokens in prompt), G. tks. (# of generated tokens), and Lat. (latency in seconds).
@@
LLM Method BEIR Covid BEIR SciFact
N@10 #Inf. P. tks. G. tks. Lat.(s) N@10 #Inf. P. tks. G. tks. Lat.(s)
NA BM25 59.5 - - - - 67.9 - - - -
Flan-T5-Large
TourRank 44.4 130.0 96946.4 1670.3 59.1 15.7 130.0 97816.3 1679.0 62.2
PRP-Graph 77.2 492.5 309267.7 - 53.7 64.6 492.0 315965.2 - 52.9
Setwise-Heapsort 75.4 129.6 58286.4 648.0 8.7 62.0 119.2 54641.1 596.0 7.9
Setwise-Insertion 74.2 120.3 53851.7 96.5 5.6 56.2 148.1 67053.6 88.2 6.7
Realm 78.4 81.2 36365.8 - 4.3 69.3 69.8 31901.7 - 3.7
BEIR DBpedia BEIR NFCorpus
NA BM25 31.8 - - - - 32.2 - - - -
Flan-T5-Large
TourRank 28.8 130.0 95231.8 1466.0 56.8 30.6 130.0 93211.3 1551.3 43.4
PRP-Graph 44.0 491.4 236478.8 - 43.3 33.9 349.2 214268.8 - 32.1
Setwise-Heapsort 41.3 124.5 41529.0 622.3 8.1 32.4 87.8 38856.5 438.3 6.5
Setwise-Insertion 42.0 109.3 36393.6 93.6 4.6 32.0 80.3 35314.2 76.4 3.9
Realm 45.9 75.4 25848.0 - 3.6 36.4 50.2 22222.4 - 2.7
Table 2: Supplementary evaluation on four BEIR datasets Covid, SciFact, DBpedia, and NFCorpus, using Flan-T5-Large model.
@@
Supplementary evaluation on BEIR datasets. We further evaluate all methods on four BEIR datasets: Covid, SciFact, DBpedia, and NFCorpus, using Flan-T5-Large model; the corresponding results are reported in Table 2. These results are consistent with our claims: Realm outperforms all existing approaches. In particular, Realm improves NDCG@10 by 2.627.62.6-27.6, while reducing the number of inferences by $39.6-84.8\%$ and cutting inference time by $30.8-93.6\%$. In terms of prompt token usage, Realm lowers cost by $39.6-89.2\%$. Note that in NFCorpus, many queries have fewer candidate passages than in other datasets, which explains the shorter processing time.
@@ 4.3 Analysis
Removing any of these components leads to a consistent drop in performance. Disabling uncertainty modeling (w/o modeling) results in a 1.53.31.5-3.3 decrease in NDCG@10 across all models, highlighting the value of Gaussian-based relevance modeling. The absence of recursive reasoning (w/o recursive) also causes noticeable degradation, underscoring the benefit of multi-round refinement. Lastly, disabling pivot-centric optimization (w/o optimization) nearly doubles latency-for example, from $7.4$s to $12.1$s with Flan-T5-XXL-confirming that our pivot selection and partitioning strategy substantially improves efficiency without compromising effectiveness.
Hyperparameter analysis. We analyze the mixing coefficient $\lambda$ introduced in Section 3.3: specifically, we sweep $\lambda\in\{0,\tfrac{1}{3},\tfrac{2}{3},1\}$ and evaluate with Flan-T5-Large, XL, XXL models on TREC DL 2020 (see Fig. 3). The results reveal a clear cost-quality trade-off; overall, $\lambda=\tfrac{2}{3}$ yields the best performance-efficiency balance, which we adopt as the default setting in our main experiments.
Figure 3: Analysis of Hyperparameter $\lambda$.
@@ 5 Conclusion
Our current design choices are partially constrained by resource and space limitations. While our relevance modeling is, in principle, compatible with a broad range of re-ranking methods, we center our evaluation on the specific framework we designed that achieved the strongest empirical performance. Besides, we have not considered skew-normal distributions for modeling relevance; while they may better capture the skewed score distributions given by the retriever, they would introduce additional complexity and inference costs. We leave the theoretical and empirical investigation of it to future work. Last, we conduct experiments using open-source models, including the Flan-T5 and LLaMA 3 series. We did not include closed-source models such as GPT-3.5 or GPT-4 accessed via API. Future extensions may consider such models to provide a more complete empirical picture of Realm.
Acknowledgments
We appreciate the anonymous reviewers for their constructive comments that helped refine and expand this work. This work was in part supported by the NSF CAREER Award No. 2326141, and NSF Awards 2212370, 2319880, 2328948, 2411294, and 2417750. This work was also performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract No. DE-AC52-07NA27344, and was supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research (ASCR) program, under SC-21. LLNL-CONF-2010514. This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. The United States Government retains, and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes.
@@ B.1 Detailed Explanation of Datasets
The TREC Deep Learning (DL) 2019 Craswell et al. (2020) and 2020 Craswell et al. (2021) datasets are benchmark collections designed to evaluate document ranking systems in complex information retrieval tasks. Both datasets are based on queries derived from real-world search logs and are built on top of the MS MARCO Bajaj et al. (2018) passage and document corpora. The TREC DL 2019 dataset includes $43$ queries with graded relevance judgments, while the 2020 version expands the test set to $54$ queries. All documents are written in English and drawn from a large web-scale corpus, with each query typically associated with hundreds to thousands of candidate passages.
We also evaluate on four datasets from the BEIR benchmark Thakur et al. (2021): Covid, SciFact, DBPedia, and NFCorpus. These collections complement TREC DL by covering diverse domains such as biomedical search, scientific fact verification, entity-centric retrieval, and consumer health information, thereby broadening the evaluation scope beyond general web search.
Together, these datasets emphasize fine-grained relevance estimation and are widely adopted for benchmarking re-ranking methods. For all experiments, the reported NDCG@10 scores are averaged over the entire dataset with a single run, ensuring stable and reliable evaluation results.